Random Point Fields, Random Measures and Bose-Einstein Condensation

HIROSHI TAMURA

Kanazawa University

The position distributions of quantum statistical mechanical models of boson gases are described by the random point fields[RPF]. We apply this old idea to obtain RPFs which describe the position distributions of simple models such as the ideal boson gases in confining potentials in Bose-Einstein condensation[BEC]. The RPFs are constructed from the elementary quantum mechanics without refering C^{∗} algebra or quantum field theories. Precise informations of the position distribution of BEC system, such as large deviation property are discussed.